OpenStudy (anonymous):

what are the 4 rules of probability?

OpenStudy (unklerhaukus):

sum of probabilities equals one

OpenStudy (anonymous):

The predictions Can be wrong..........lol

OpenStudy (anonymous):

I think it should be probability should not exceed 1

OpenStudy (anonymous):

Because sum can be any

OpenStudy (anonymous):

There is the (And and Or) Rule

OpenStudy (anonymous):

It happens or else it doesn't. The probabilty of an event happening added the probability of it not happing is always 1. P(A happens) + P(A doen't happen) = 1 2. Exclusivity. If A and B can't both happen at the same time (in which case we say that A and B are mutually exclusive), then P(either A or B happens) = P(A happens) + P(B happens) 3. Independence. If B is no more or less likely to happen when A happens than when A doesn't (in which case we say that A and B are independent), then P(A and B both happen) = P(A happens) * P(B happens) 4. Sub-Events. If whenever A happens B must also happen, then B must be at least as likely as A, so P(A happens) <= P(B happens)

OpenStudy (anonymous):

Now in the And rule u multiply and in the Or rule u add

OpenStudy (anonymous):

This question is inexact (and who says there are only 4 rules anyway.-)

OpenStudy (unklerhaukus):

axioms ?

OpenStudy (anonymous):

Should be like that, there is more than one version of probability theory out there. And it seems this question is only addressing the theory not the practice (stats)

OpenStudy (anonymous):

You can look up Skyrmes SIX rules of probability.